• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 16:07:42 on October27,2001

    Entry number 75205

    keyword=shift summary

    white spectra scan was completed according to the run plan
    Left HRS was put at 4.3 GeV with sieve slit for RCS optics run
    Q1 power trip one time. Scott Spiegel reset it.